# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v2

@@ 4.4 Determination of the graph structure
A crucial issue when dealing with spatio-temporal GNNs for time series problems is the determination of the graph structure, i.e., the connectivity of the nodes. Namely, some time series datasets come by their own nature with a pre-defined graph structure (e.g., a road network), whereas some others do not. When a natural pre-defined graph structure is available, it helps the model to better capture the underlying dynamics of the system. When it is not directly available, it must be defined by the user in some way (based on domain knowledge or some metrics), or determined by some algorithm (such as visibility graphs [visibility_Gori, visibility_graph, art_14]), or learned by the model itself.
@@
Pre-defined graphs are typically generated based on prior knowledge and specific metrics. These methods for the construction of the graphs typically rely on distance or similarity information. As for location-based techniques, they utilize geographical coordinates (latitude and the longitude) of the nodes to calculate the distance between them (e.g., using haversine or geodesic distance). This distance information is then used to determine the edge weights with techniques such as the thresholded Gaussian kernel (which applies a negative exponential function to the distance), weighted $k$-NN, or reversed MinMax. As for similarity-based techniques instead, they rely on similarity (or dissimilarity) of signal data. These techniques include methods such as correlation, cosine similarity, dynamic time warping (DTW), maximum information coefficient (MIM), or transfer entropy. In general, no single approach stands out as a clear winner, and performance should be evaluated on each dataset. For references and a more detailed explanation of these and other techniques for constructing pre-defined structures, see Ref. [art_160].
@@
As for automatic graph structure learning techniques, a distinction can be made between models that rely on predefined graph structures (sometimes referred to as supplementary graph structure learning models), and those that learn the graph structure entirely from scratch. In the first category, there are models such as Graph WaveNet (GWN) [GWN], which constructs an adaptive matrix as a complement to a predefined graph structure [art_405]. These techniques still rely on some form of prior assumptions. To address this limitation, models that learn the graph structure from scratch have been proposed. One of them is the MTGNN model [MTGNN], which extracts uni-directional relationships from node embeddings as it learns, and achieves state-of-the-art results across different datasets. It is one of the first models to use self-learned graph structures [art_417]. When the graph structure is learned, some authors introduce a graph regularization term to the training loss function to optimize the learned graph adjacency matrix. This regularization term is typically designed to control the sparsity of the matrix, and to promote graph proximity, encouraging larger entries when the series in the corresponding nodes are “close enough” [art_448]. Some researchers argue that models which generate graphs without prior knowledge may overlook local conditions when constructing adjacency matrices [art_397]. Therefore, it is important to conduct further research in this area to identify the most effective techniques for generating informative graphs. For a more detailed explanation of graph learning models, see Refs. [art_396, art_405].
@@
Another key aspect in the determination of the graph structure is the choice between using a static or dynamic adjacency matrix, i.e. whether to use the same matrix for all graphs or not. When working with graphs defined in advance by specific metrics, it is common to use either a static matrix or a dynamic matrix generated through a sliding window approach. In contrast, when graphs are learned by the model, dynamic adjacency matrices are typically employed. The use of a dynamic adjacency matrix has the advantage of capturing the dynamic nature of the problem [art_447]. However, the learned dynamic graphs often show significant differences between adjacent time steps, whereas real-world graph structures typically evolve smoothly over time. Therefore, this approach may not accurately capture the stable evolution observed in the real world.
@@
Consequently, some researchers are exploring an intermediate approach, assuming that adjacent graph structures follow a basic pattern. Instead of defining static or dynamic adjacency matrices, they fuse a static backbone graph with a dynamic adjacency matrix. The static network captures the common graph structure over time, while the dynamic network focuses on learning the changes over time. An example of this intermediate approach can be found in Ref. [art_415], which also provides a comparison between this intermediate approach and the classical ones.
@@
In general, there is no rule that can determine in advance the best approach to follow. In fact, it depends on both the application domain and the specific dataset. For this reason, we do not explore this aspect further. Each subsection of the thematic analysis in Sec. 5 discusses the most prevalent approach for each group of papers.
@@ 4.5 Model evaluation
Evaluating the performance of a time series forecasting or classification model is essential to assess its reliability. Moreover, experimental results often show that even when datasets involve the same forecasting task, the transferability of model architectures is limited [art_408]. Therefore, a proper evaluation of the model is a crucial aspect. A first key aspect to consider is its accuracy. The accuracy of a model can be measured by many different metrics, which vary between classification and forecasting tasks. Each metric provides its own insight into different aspects of prediction quality, so multiple metrics are typically used for a more comprehensive evaluation. In addition to presenting a table of the error metrics computed for the entire graph, some authors provide a spatial visualization of these errors across the various nodes, as for example in Refs. [art_28, art_449].
@@
As for classification problems, common evaluation metrics include accuracy, precision, recall, and F1 score. Given $TP$, $TN$, $FP$, and $FN$ the number of true positives, true negatives, false positives, and false negatives, respectively, the formulas for the binary classification problem are reported in Tab. 6.
@@
Similar formulas can be applied for the multi-class classification problem.
Besides accuracy, another important aspect to consider is the complexity of the model. Finding an optimal balance between accuracy and complexity is crucial for developing models that are not only accurate but also efficient. For this reason, some authors include the training and inference computational times, as well as the number of parameters of their models, allowing for a comparison of these aspects across different models. However, not all authors do so, and even if they do it, the characteristics of the hardware and other settings used by each author must be taken into account when comparing computational times. Consequently, it was not possible to provide an accurate and detailed comparison of the relationship between complexity and performance across the different models in this review. Although creating a plot to compare accuracy and computational times of the different models would be very useful (as in Ref. [art_407]), it is not possible to provide such a comparison here. The authors of the various papers should be encouraged to incorporate such comparisons in their work.
@@ 5 Thematic analysis
In each following subsection, one per theme, a list of the benchmark models that appear in the selected journal papers will be provided. Within each theme, they will be classified as: mathematical and statistical models coming from Econometrics, non-GNN machine learning models, and GNN-based models. Mathematical and statistical models are characterized by the fact that they generate forecasts by analyzing a limited amount of data, and they usually struggle with cross-sectional large-scale and temporal long-term dependencies. Non-GNN machine learning models use more complex and highly nonlinear architectures, mainly neural networks, but they are not necessarily able to capture intricate spatio-temporal dependencies within the data. GNN-based models, conversely, attempt to represent these complex dependencies using various techniques, resulting in better, though not perfect, accuracy. In addition, information on the most common datasets will be given, along with the best results reported in the literature. All results presented in the tables of this SLR are taken directly from the tables of the selected papers, as we did not conduct any further experiments to obtain additional results. Note that in the tables which list benchmark models we have included references to any GNN model found. In the results tables, instead, we have explicitly referenced only those models proposed in the selected papers, in order to highlight them. Further information about the models used as benchmarks can be found either in the benchmark tables or in the papers where they are used as benchmarks.
@@ 5.1 Energy
The first thematic group is “Energy” (with 33 out of 281 papers), which includes all the studies related to energy systems. In particular, it includes research on electricity load, energy consumption, and power generation forecasting. These topics are somehow connected to those of the “Environment” subsection discussed later (Subsec. 5.2). For example, green energy sources are associated with environmental phenomena such as wind speed and solar radiation inflow. Our rationale for this selection is that here we focus on energy applications, whereas Subsec. 5.2 will refer to the study of environmental data that may or may not be related to energy applications. Wind speed forecasting, for instance, is connected to wind power generation due to the cubic relationship between wind speed and wind power [art_29]. However, wind speed forecasting may also be of interest for other reasons, and thus it is included in the more general Subsec. 5.2.
@@ 5.1.1 Overview
An overview of the selected papers reveals a relevant interest in the application of GNNs in the energy field starting from 2022. The most frequently studied fields include forecasting of wind [art_20, art_24, art_27, art_332, art_333] and solar power [art_19, art_21, art_22, art_26, art_349], which are renewable energy sources characterized by their intermittency. The most popular journal among the selected papers is Applied Energy by Elsevier, with 5 published papers. The majority of the corresponding authors are affiliated with institutions in China.
@@ 5.1.2 Datasets
Commercial and Residential Load [art_348] https://www.osti.gov/biblio/1788456
Danish smart heat meters [art_15] https://doi.org/10.5281/zenodo.6563114
DKA Solar Centre [art_19] https://dkasolarcentre.com.au/
IEEE bus systems [art_23] https://doi.org/10.1109/UPEC.2015.7339813 (source paper)
Individual Household Electric Power Consumption [art_344] https://archive.ics.uci.edu/ml/datasets/individual+household+electric+power+consumption
JERICHO-E-usage [art_15] https://doi.org/10.6084/m9.figshare.c.5245457.v1
NREL [art_20], [art_22], [art_24] https://www.nrel.gov/
OPSD [art_16] https://open-power-system-data.org/
PJM Hourly Energy Consumption [art_390] https://www.pjm.com/markets-and-operations/data-dictionary
PVGIS [art_349] http://re.jrc.ec.europa.eu/pvgis/apps4/pvest.php
PV Switzerland [art_21], [art_22] https://doi.org/10.1109/IJCNN48605.2020.9207573 (source paper)
Residential heat and electricity demand [art_341] https://ieee-dataport.org/open-access/8-years-hourly-heat-and-electricity-demand-residential-building
SGCC dataset [art_342] http://www.sgcc.com.cn/
Southern Power Grid [art_344] https://www.csg.cn/
UMass Smart* Dataset [art_17] https://traces.cs.umass.edu/docs/traces/smartstar/
Three of the selected papers [art_20, art_22, art_24] use data from the National Renewable Energy Laboratory (NREL), a national laboratory of the U.S. Department of Energy (https://www.nrel.gov/). Additional datasets were used for exogenous weather variables and satellite imagery (such as Himawari-8, available at https://www.data.jma.go.jp/mscweb/en/himawari89/space_segment/spsg_sample.html in Ref. [art_19]), but they are of course highly dependent on the location of the study. The data time granularity ranges from 1 minute to 1 day, and the forecasting horizon from tens of minutes to days. These quantities depend on the dataset, the domain of application, and the purpose of the forecast. Most datasets require pre-processing, including missing data interpolation, normalization, and satellite image processing.
@@ 5.1.3 Types of models
As for the taxonomy of the proposed GNN models, 17 out of 33 papers use a convolutional GNN, and the other 9 use an attentional GNN. Six papers from 2024 use both convolutional and attentional approaches, and one paper [art_338] uses the original recurrent approach on a graph defined via the visibility algorithm. As to the graph structure used, the papers typically provide only a brief description. For example, only few papers highlight whether the graph structure is static or dynamic, and the way in which it is constructed is often described summarily. In some papers, the graph structure reflects the geographic location of the elements (e.g., wind or solar power plants, as in Refs. [art_24, art_26, art_349]), whereas in others it is based on the correlation or similarity between time series (as in Ref. [art_19]). In some studies, a pre-defined structure is imposed and the weighted adjacency matrix is learned directly from the model, often using the attention mechanism as in Ref. [art_22]. The majority of papers in this group rely on a static graph.
The most common loss function used to train these models is the MSE. Many papers mention the use of the programming languages Python and Matlab, and few of them mention the use of the Python libraries TensorFlow [2], PyTorch [PyTorch], and PyTorch Geometric [PyTorch_Geometric] for the implementation of the GNN model. In this group, only Ref. [art_17] provides the link to the source code of the proposed residential load forecasting with multiple correlation temporal graph neural networks (RLF-MGNN) model (available at https://codeocean.com/capsule/9294192/tree/v1).
@@ 5.1.4 Benchmark models
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [art_15], [art_16], [art_22], [art_25], [art_333], [art_337], [art_338], [art_341], [art_343], [art_349]
Historical average (HA) [art_350], [art_390]
Vector ARIMA [art_332], [art_333], [art_343], [art_350], [art_390]
Non-GNN machine learning methods Autoformer [Autoformer] [art_333], [art_341]
Bidirectional long-short term memory network (BiLSTM) [BiLSTM] [art_15], [art_19], [art_339]
Convolutional neural network (CNN) [art_15], [art_18], [art_21], [art_22], [art_343], [art_347]
Hybrid convolutional neural network and long-short term memory network (CNN-LSTM) [art_15], [art_17], [art_351]
Encoder-decorer architectures (ED) [art_15], [art_19], [art_21], [art_22]
Feed-forward neural network (FNN) [art_15], [art_18], [art_19], [art_25], [art_27], [art_351]
Gated recurrent unit (GRU) [art_15], [art_25], [art_334], [art_338], [art_343], [art_347]
Informer [Informer] [art_15], [art_16]
$k$-nearest neighbors (KNN) [art_18], [art_24], [art_342], [art_343]
Long-short term memory network (LSTM) [art_15], [art_16], [art_19], [art_20], [art_23], [art_24], [art_333], [art_334], [art_335], [art_340], [art_341], [art_343], [art_344], [art_347], [art_348], [art_349], [art_353]
Long-short term time series network (LSTNet) [LSTNet] [art_15], [art_332], [art_334]
Recurrent neural network (RNN) [art_16], [art_17], [art_19], [art_20]
Seq2Seq [Seq2seq] [art_335], [art_343]
Support vector regression (SVR) [art_21], [art_22], [art_24], [art_27], [art_338], [art_342], [art_343], [art_344], [art_348]
Transformer [attention] [art_16], [art_19], [art_335], [art_343]
Extreme gradient boosting (XGBoost) [art_25], [art_334], [art_335], [art_341]
GNN methods Diffusion convolutional recurrent neural network (DCRNN) [DCRNN] [art_349], [art_350], [art_390]
Multivariate time series forecasting with graph neural network (MTGNN) [MTGNN] [art_332], [art_340], [art_341]
Spatio-temporal autoregressive model (ST-AR) [ST-AR] [art_21], [art_22]
Spatio-temporal graph convolutional network (STGCN) [STGCN] [art_339], [art_341], [art_348], [art_351], [art_353]

### v4

@@ 4.4 Determination of the graph structure
A crucial issue when dealing with spatio-temporal GNNs for time series tasks is the determination of the graph structure, i.e., the connectivity of the nodes. The type of spatial information captured by the graph can vary widely across domains. For instance, in traffic networks, nodes often represent sensors or intersections, and edges correspond to actual roads or physical connections. In contrast, in domains such as healthcare or finance, nodes may represent patients, brain regions, or financial instruments, and edges encode correlations, interactions, or other relational information. This variability across datasets highlights the importance of carefully designing the graph structure to reflect the spatial relationships relevant to each domain. The specific choices made in different application areas are discussed in the corresponding subsections of this review.
Namely, some time series datasets come by their own nature with a pre-defined graph structure (e.g., a road network for traffic datasets), whereas some others do not. When a natural graph structure is available, it helps the model capture the underlying dynamics more effectively. When it is not directly available, it must be defined by the user in some way (based on domain knowledge or some metrics), or determined by some algorithm (such as visibility graphs [468, 201, 383]), or learned by the model itself.
@@
Pre-defined graphs are typically generated based on prior knowledge and specific metrics. These methods for the construction of the graphs typically rely on distance or similarity information. As for location-based techniques, they utilize geographical coordinates (latitude and the longitude) of the nodes to calculate the distance between them (e.g., using haversine or geodesic distance). This distance information is then used to determine the edge weights with techniques such as the thresholded Gaussian kernel (which applies a negative exponential function to the distance), weighted $k$-NN, or reversed MinMax. As for similarity-based techniques instead, they rely on similarity (or dissimilarity) of signal data. These techniques include methods such as correlation, cosine similarity, dynamic time warping (DTW), maximum information coefficient (MIM), or transfer entropy. In general, no single approach stands out as a clear winner, and performance should be evaluated on each dataset. For references and a more detailed explanation of these and other techniques for constructing pre-defined structures, see Ref. [15].
As for automatic graph structure learning techniques, a distinction can be made between models that rely on pre-defined graph structures (sometimes referred to as supplementary graph structure learning models), and those that learn the graph structure entirely from scratch. Two notable examples from Wu et al. illustrate these approaches.
The first example is Graph WaveNet (GWN) [425], which constructs a self-adaptive adjacency matrix $A$ using two randomly initialized learnable node embedding matrices $E_{1}$ and $E_{2}$ as:
$A=\text{softmax}\left(\text{ReLU}\left(E_{1}\,E_{2}^{\top}\right)\right)\>,$ (5)
where $E_{1}$ and $E_{2}$ correspond to the source and target node embeddings, respectively. Their multiplication represents the spatial dependency weights between source nodes and target nodes. In the original work, these learned hidden graph dependencies are combined with a pre-defined graph structure. The authors note that when such a structure is unavailable, the self-adaptive adjacency matrix alone can be used to model hidden spatial dependencies.
As a second example, which overcomes the reliance on a pre-defined graph, the same authors proposed the MTGNN model [424], which extracts uni-directional relationships from node embeddings as follows:
$M_{1}=\tanh(\eta\,E_{1}\Phi_{1})$ (6)
$M_{2}=\tanh(\eta\,E_{2}\Phi_{2})$
$A=\text{ReLU}\left(\tanh\left(\eta\left(M_{1}M_{2}^{\top}-M_{2}M_{1}^{\top}\right)\right)\right)$
$\text{idx}=\text{argtopk}(A[i,:])\quad\forall i=1,\ldots,n$
$A[i,-\text{idx}]=0\>,$
where $E_{1}$ and $E_{2}$ are randomly initialized learnable node embedding matrices, $\Phi_{1}$ and $\Phi_{2}$ are model parameters, $\eta$ is a hyperparameter controlling the saturation rate of the activation function, and $\text{argtopk}(\cdot)$ is a function that returns the index of the top-k largest values of a vector. Unlike GWN, which can be used either standalone or to augment a pre-defined graph, MTGNN learns the graph structure fully end-to-end. These two models are among the first to explore self-learned graph structures. Although other methods have since been proposed, GWN and MTGNN continue to serve as prototypical models for self-learned graph structures.
@@
When the graph structure is learned, some authors introduce a graph regularization term to the training loss function to optimize the learned graph adjacency matrix. This regularization term is typically designed to control the sparsity of the matrix, and to promote graph proximity, encouraging larger entries when the series in the corresponding nodes are “close enough”. Some researchers argue that models which generate graphs without prior knowledge may overlook local conditions when constructing adjacency matrices [152]. Therefore, it is important to conduct further research in this area to identify the most effective techniques for generating informative graphs. For a more detailed explanation of graph learning models, see Refs. [137, 240].
@@
Another key aspect in the determination of the graph structure is the choice between using a static or dynamic adjacency matrix, i.e. whether to use the same matrix for all graphs or not. When working with graphs defined in advance by specific metrics, it is common to use either a static matrix or a dynamic matrix generated through a sliding window approach. In contrast, when graphs are learned by the model, dynamic adjacency matrices are typically employed. The use of a dynamic adjacency matrix has the advantage of capturing the dynamic nature of the problem [126]. However, the learned dynamic graphs often show significant differences between adjacent time steps, whereas real-world graph structures typically evolve smoothly over time. Therefore, this approach may not accurately capture the stable evolution observed in the real world.
@@
Consequently, some researchers are exploring an intermediate approach, assuming that adjacent graph structures follow a basic pattern. Instead of defining static or dynamic adjacency matrices, they fuse a static backbone graph with a dynamic adjacency matrix. The static network captures the common graph structure over time, while the dynamic network focuses on learning the changes over time. An example of this intermediate approach can be found in Ref. [285], which also provides a comparison between this intermediate approach and the classical ones.
@@
In general, there is no rule that can determine in advance the best approach to follow. In fact, it depends on both the application domain and the specific dataset. For this reason, we do not explore this aspect further. As an example, for a detailed discussion of graph construction techniques in traffic flow prediction models, the reader is referred to Ref. [78]. Each subsection of the thematic analysis in Sec. 5 discusses the most prevalent approach for each group of papers.
@@ 4.5 Model evaluation
Evaluating the performance of a time series forecasting or classification model is essential to assess its reliability. Moreover, experimental results often show that even when datasets involve the same forecasting task, the transferability of model architectures is limited [32]. Therefore, a proper evaluation of the model is a crucial aspect. A first key aspect to consider is its accuracy. The accuracy of a model can be measured by many different metrics, which vary between classification and forecasting tasks. Each metric provides its own insight into different aspects of prediction quality, so multiple metrics are typically used for a more comprehensive evaluation. In addition to presenting a table of the error metrics computed for the entire graph, some authors provide a spatial visualization of these errors across the various nodes, as for example in Ref. [437].
@@
As for classification problems, common evaluation metrics include accuracy, precision, recall, and F1 score. Given $TP$, TNTN, FPFP, and $FN$ the number of true positives, true negatives, false positives, and false negatives, respectively, the formulas for the binary classification problem are reported in Tab. 6. Similar formulas can be applied for the multi-class classification problem.
@@
Besides accuracy, another important aspect to consider is the complexity of the model. Finding an optimal balance between accuracy and complexity is crucial for developing models that are not only accurate but also efficient. For this reason, some authors include the training and inference computational times, as well as the number of parameters of their models, allowing for a comparison of these aspects across different models. However, not all authors do so, and even if they do it, the characteristics of the hardware and other settings used by each author must be taken into account when comparing computational times. Consequently, it was not possible to provide an accurate and detailed comparison of the relationship between complexity and performance across the different models in this review. Although creating a plot to compare accuracy and computational times of the different models would be very useful (as in Refs. [140] and [81]), it is not possible to provide such a comparison here. The authors of the various papers should be encouraged to incorporate such comparisons in their work.
@@ 5 Thematic analysis
In each following subsection, one per theme, a list of the benchmark models that appear in the selected papers will be provided. Within each theme, they will be classified as: mathematical and statistical models coming from Econometrics, non-GNN machine learning models, and GNN-based models. Mathematical and statistical models are characterized by the fact that they generate forecasts by analyzing a limited amount of data, and they usually struggle with cross-sectional large-scale and temporal long-term dependencies. Non-GNN machine learning models use more complex and highly nonlinear architectures, mainly neural networks, but they are not necessarily able to capture intricate spatio-temporal dependencies within the data. GNN-based models, conversely, attempt to represent these complex dependencies using various techniques, resulting in better, though not perfect, accuracy. In addition, information on the most common datasets will be given, along with the best results reported in the literature. All results presented in the tables of this SLR are taken directly from the tables of the selected papers, as we did not conduct any further experiments to obtain additional results. When examining the results tables, the reader is invited to consult the GitHub repository associated with this review, which provides interactive functionality for sorting and filtering models by any metric and generating plots, making it easy to analyze and compare the results. Note that in the tables which list benchmark models we have included references to any GNN model found. In the results tables, instead, we have explicitly referenced only those models proposed in the selected papers, in order to highlight them. Further information about the models used as benchmarks can be found either in the benchmark tables or in the papers where they are used as benchmarks. The most common benchmarks used across all groups of papers are also listed in the GitHub repository.
@@ 5.1 Energy
The first thematic group is “Energy” (with 33 out of 366 papers), which includes all the studies related to energy systems. In particular, it includes research on electricity load, energy consumption, and power generation forecasting. These topics are somehow connected to those of the “Environment” subsection discussed later (Subsec. 5.2). For example, green energy sources are associated with environmental phenomena such as wind speed and solar radiation inflow. Our rationale for this selection is that here we focus on energy applications, whereas Subsec. 5.2 will refer to the study of environmental data that may or may not be related to energy applications. Wind speed forecasting, for instance, is connected to wind power generation due to the cubic relationship between wind speed and wind power [11]. However, wind speed forecasting may also be of interest for other reasons, and thus it is included in the more general Subsec. 5.2.
@@ 5.1.1 Overview
An overview of the selected papers reveals a relevant interest in the application of GNNs in the energy field starting from 2022. The most frequently studied fields include forecasting of wind [132, 460, 216, 439, 452] and solar power [350, 351, 317, 377], which are renewable energy sources characterized by their intermittency. The most popular journal among the selected papers is Applied Energy by Elsevier, with 5 published papers, and the only “Energy” conference paper was presented at IJCAI.
@@ 5.1.2 Datasets
Commercial and residential load [87] https://www.osti.gov/biblio/1788456
Danish smart heat meters [405] https://doi.org/10.5281/zenodo.6563114
IEEE bus systems [509] https://doi.org/10.1109/UPEC.2015.7339813 (source paper)
Individual household electric power consumption [156] https://archive.ics.uci.edu/ml/datasets/individual+household+electric+power+consumption
JERICHO-E-usage [405] https://doi.org/10.6084/m9.figshare.c.5245457.v1
NREL [132], [351], [460] https://www.nrel.gov/
OpenEI residential load [336] https://openei.org/datasets/files/961/pub/RESIDENTIAL_LOAD_DATA_E_PLUS_OUTPUT/HIGH/
OPSD [486] https://open-power-system-data.org/
PJM hourly energy consumption [319] https://www.pjm.com/markets-and-operations/data-dictionary
PVGIS [377] http://re.jrc.ec.europa.eu/pvgis/apps4/pvest.php
PV Switzerland [350], [351] https://doi.org/10.1109/IJCNN48605.2020.9207573 (source paper)
Residential heat and electricity demand [288] https://ieee-dataport.org/open-access/8-years-hourly-heat-and-electricity-demand-residential-building
SGCC dataset [161] http://www.sgcc.com.cn/
Southern power grid [156] https://www.csg.cn/
UMass smart* [398] https://traces.cs.umass.edu/docs/traces/smartstar/
Three of the selected papers [132, 351, 460] use data from the National Renewable Energy Laboratory (NREL), a national laboratory of the U.S. Department of Energy (https://www.nrel.gov/). The data time granularity ranges from 1 minute to 1 day, and the forecasting horizon from tens of minutes to days. These quantities depend on the dataset, the domain of application, and the purpose of the forecast. Most datasets require pre-processing, including missing data interpolation, normalization, and satellite image processing.
@@ 5.1.3 Types of models
As for the taxonomy of the proposed GNN models, 17 out of 33 papers use a convolutional GNN, and 8 use an attentional GNN. Seven papers use both convolutional and attentional approaches, and one paper [104] uses the original recurrent approach on a graph defined via the visibility algorithm. As to the graph structure used, the papers typically provide only a brief description. For example, only few papers highlight whether the graph structure is static or dynamic, and the way in which it is constructed is often described summarily. In some papers, the graph structure reflects the geographic location of the elements (e.g., wind or solar power plants, as in Refs. [460, 317, 377]), whereas in others it is based on the correlation or similarity between time series (as in Ref. [323]). In some studies, a pre-defined structure is imposed and the weighted adjacency matrix is learned directly from the model, often using the attention mechanism as in Ref. [351]. The majority of papers in this group rely on a static graph.
The most common loss function used to train these models is the MSE. Many papers mention the use of the programming languages Python and Matlab, and few of them mention the use of the Python libraries TensorFlow [2], PyTorch [309], and PyTorch Geometric [91] for the implementation of the GNN model. In this group, only Ref. [398] provides the link to the source code of the proposed residential load forecasting with multiple correlation temporal graph neural networks (RLF-MGNN) model (available at https://codeocean.com/capsule/9294192/tree/v1).
@@ 5.1.4 Benchmark models
Mathematical and statistical methods Autoregressive integrated moving average (ARIMA) [405], [486], [351], [147], [452], [323], [104], [288], [324], [377], [256]
Historical average (HA) [500], [319]
Vector ARIMA [439], [452], [324], [500], [319]
Non-GNN machine learning methods Autoformer [414] [452], [288]
Bidirectional long-short term memory network (BiLSTM) [109] [405], [477]
Convolutional neural network (CNN) [405], [93], [350], [351], [324], [24]
Hybrid convolutional neural network and long-short term memory network (CNN-LSTM) [405], [398], [315]
Encoder-decoder architectures (ED) [405], [350], [351]
Feed-forward neural network (FNN) [405], [93], [147], [216], [315], [256]
Gated recurrent unit (GRU) [405], [147], [531], [104], [324], [24]
Informer [520] [405], [486]
$k$-nearest neighbors (KNN) [93], [460], [161], [324]
Long-short term memory network (LSTM) [405], [486], [132], [509], [460], [452], [531], [148], [265], [288], [324], [156], [24], [87], [377], [463], [256]
Long-short term time series network (LSTNet) [202] [405], [439], [531]
Recurrent neural network (RNN) [486], [398], [132]
Seq2Seq [365] [148], [324]
Support vector regression (SVR) [350], [351], [460], [216], [104], [161], [324], [156], [87], [256]
Transformer [375] [486], [148], [324]
Extreme gradient boosting (XGBoost) [147], [531], [148], [288]
GNN methods Diffusion convolutional recurrent neural network (DCRNN) [237] [377], [500], [319]
Multivariate time series forecasting with graph neural network (MTGNN) [424] [439], [265], [288]
Spatio-temporal autoregressive model (ST-AR) [29] [350], [351]
Spatio-temporal graph convolutional network (STGCN) [457] [477], [288], [87], [315], [463]